OpenStudy (anonymous):

A go cart racetrack has 100-meter straightways and semicircular ends with diameter of 40 meters. Calculate the average speed in meters per minute of a go cart if it completes 4 laps in 6 minutes.

OpenStudy (anonymous):

the length of circumference for a semicircle is 1/2 *pi *diameter. =1/2 * pi * 40 = 20pi. the length of a lap is 100+20pi. to find the average speed, divide the total distance by total time. in your case, average speed= 4 laps / 6 minutes.

OpenStudy (anonymous):

thanks!